# Direct CKM determination from W decays at future lepton colliders

### v1

@@ 6 Conclusions
In this work we showed the potential of FCC-ee to measure $|V_{cs}|$ and $|V_{cb}|$ from hadronic $W$ decays, with similar results expected at other future lepton colliders. Our findings are summarized in Table 1. In both cases, the relative uncertainty will largely surpass present levels. Assuming per-mille level systematics on the tagger parameters, one can expect to increase the precision on CKM elements by one order of magnitude. More conservative assumptions for the tagger performances will lead to smaller improvement, especially in the systematic dominated case of $|V_{cs}|$. Together with the systematic level, final sensitivities strongly depend on the rejection capability of the taggers, that is on the mistag probabilities. Projections on tagger working points at FCC-ee indicate that such low mistag rates can be achieved; particular attention will be needed for $s$-tagger, where large $s$-to-light jet mistagging can sizeably degrade the results.
@@ Acknowledgments
The authors would like to thank Jernej Kamenik, Stephane Monteil, Marie-Hélène Schune, Michele Selvaggi and Jure Zupan for useful discussions. MS acknowledges support in part by the DOE grant de-sc0011784 and NSF OAC-2103889. DM acknowledges support by the MUR grant PRIN 20224JR28W.
@@ Appendix A Details on the probability function
where ${\cal B}_{ij}$ are the $W$ branching ratios in each channel defined in Eq. (3) and ${\cal A}_{W}$ is the acceptance. In can be explicitly checked that the sum of these 6 probabilities gives precisely ${\cal B}_{\rm had}{\cal A}_{W}$, as expected. Finally, the expected number of events $N_{B_{bc;1},B_{bc;2}}$ is obtained as

### v2

@@ 6 Conclusions
In this work we showed the potential of FCC-ee, or CEPC, to measure $|V_{cs}|$ and $|V_{cb}|$ from hadronic $W$ decays. [^3] Our results could also be applied to ILC by choosing a different working point for the tagging efficiencies in Fig. 3. Our findings are summarized in Table 1. In both cases, the relative uncertainty will largely surpass present levels. Assuming per-mille level systematics on the tagger parameters, one can expect to increase the precision on CKM elements by one order of magnitude. More conservative assumptions for the tagger performances will lead to smaller improvement, especially in the systematic dominated case of $|V_{cs}|$. Together with the systematic level, final sensitivities strongly depend on the rejection capability of the taggers, that is on the mistag probabilities. Projections on tagger working points at FCC-ee indicate that such low mistag rates can be achieved; particular attention will be needed for $s$-tagger, where large $s$-to-light jet mistagging can sizeably degrade the results.
@@ Acknowledgments
The authors would like to thank Jernej Kamenik, LingFeng Li, Stephane Monteil, Marie-Hélène Schune, Michele Selvaggi and Jure Zupan for useful discussions. MS acknowledges support in part by the DOE grant de-sc0011784 and NSF OAC-2103889. DM acknowledges support by the MUR grant PRIN 20224JR28W.
@@ Appendix A Details on the probability function
where ${\cal B}_{ij}$ are the $W$ branching ratios in each channel defined in Eq. (3) and ${\cal A}_{W}$ is the acceptance. It can be explicitly checked that the sum of these 6 probabilities gives precisely ${\cal B}_{\rm had}{\cal A}_{W}$, as expected. Finally, the expected number of events $N_{B_{bc;1},B_{bc;2}}$ is obtained as